You have a rectangle that the perimeter is 34ft. The length is the width plus 9. What is the length and width?
Answer by oberobic(2304)

P = 2(L+W)
P = 34
.
L = W+9
.
34 = 2(W+9 +W)
34 = 2W +18 +2W
16 = 4W
4W = 16
W = 4
.
L = 4+9
L = 13
.
P = 2(4+13)
P = 2(17)
P = 34
.
Answer: Length = 13 ft; width = 4 ft.
.
Done.